Solve the problem from BM 13901, finding two squares, one of which has a side five units longer than the other and whose total area is 25 , 25 square units.

side note: 25,25 converted from sexagesimal is 1525
Answer #1

let square length re. (4+5) (22+5) unit? To fal Area = 25 04+5) x (x+5) = 25 (1+5)2= 25 xts = 15 2=5-5 n=-5-5 3= -10 Til = 10please upvote my answer thanks please upvote it